### v1

# Mokav: Execution-driven Differential Testing with LLMs

Abstract.
It is essential to detect functional differences in various software engineering tasks, such as automated program repair, mutation testing, and code refactoring. The problem of detecting functional differences between two programs can be reduced to searching for a difference exposing test (DET): a test input that results in different outputs on the subject programs. In this paper, we propose Mokav, a novel execution-driven tool that leverages LLMs to generate DETs. Mokav takes two versions of a program (P and Q) and an example test input. When successful, Mokav generates a valid DET, a test input that leads to different outputs on P and Q. Mokav iteratively prompts an LLM with a specialized prompt to generate new test inputs. At each iteration, Mokav provides execution-based feedback regarding previously generated tests until the LLM produces a DET. We evaluate Mokav on 1,535 pairs of Python programs collected from the Codeforces competition platform and 32 pairs of programs from the QuixBugs dataset. Our experiments show that Mokav outperforms the state-of-the-art, Pynguin and Differential Prompting (ASE 2023), by a large margin. Mokav can generate DETs for 81.7% (1,255/1,535) of the program pairs in our benchmark (versus 4.9% for Pynguin and 37.3% for Differential Prompting). We demonstrate that all components in our system, including the iterative and execution-driven approaches, contribute to its high effectiveness.
Test Generation, Large Language Models, Behavioral Difference
copyright: none
1. Introduction
In differential testing, two versions of a program are given (P & Q), and a test input is searched for such that P & Q produce different outputs (Gulzar et al., 2019). This test exposes a functional difference between P & Q, thus is called a difference exposing test (DET). DETs are useful in various software development tasks, such as for explaining changes (Castellano et al., 2022), detecting anomalies and bugs (Jarman et al., 2020a; Gulzar et al., 2019), analyzing refactorings (Daniel et al., 2007), etc. Generating a DET is significantly more challenging than random test generation because it requires exploring the vast input space of P & Q for the rare inputs that trigger the functional differences (McKeeman, 1998).
Existing techniques to find difference exposing tests limit the search scope using symbolic execution (Rutledge and Orso, 2022), type aware mutation (Liu et al., 2024c), and code coverage optimization with genetic algorithms (Da Silva et al., 2020). These techniques do not leverage recent advanced methods for understanding P & Q’s semantics to guide the search for DETs. Specifically, large language models (LLMs) have shown strong performance in program comprehension (Yuan et al., 2023). This makes LLMs an interesting candidate for difference exposing test search. Previous studies have employed LLMs to conduct various software testing tasks, such as test completion (Nie et al., 2023), test input generation (Xia et al., 2024), test oracle generation (Dinella et al., 2022), unit test generation (Schäfer et al., 2023), and GUI testing (Liu et al., 2024a). To the best of our knowledge, there is no prior work on leveraging LLMs for differential testing.
In this paper, we propose Mokav, a novel LLM-based tool for differential testing. Mokav iteratively requests an LLM to generate a DET based on two programs P & Q. At each iteration, Mokav provides three pieces of information in its prompt to the LLM. First, an example test input that produces the same output on P & Q. This example test hints at the model regarding the type and structure of inputs acceptable by P & Q. Second, Mokav runs the example test on P & Q and adds their identical output to the prompt. The output produced by P & Q on the example test provides extra information about their functionality. Third, Mokav collects variable values while running the example test on P & Q and adds detected disparities in the prompt. Using this fine-grained execution data, Mokav steers the LLM towards parts of the input space that contain DETs.
We evaluate Mokav on two datasets: 32 pairs of Python programs in the widely used QuixBugs (Lin et al., 2017) dataset, and our own crafted dataset C4DET. We create C4DET by carefully selecting 1,535 pairs of Python programs from Codeforces [^1] https://codeforces.com/ competitions, with a guaranteed functional difference. Our evaluation results show that Mokav outperforms our two strong baselines, Pynguin (Lukasczyk and Fraser, 2022) and Differential Prompting (DPP(Li et al., 2023). Mokav generates a DET for 81.7% (1,255/1,535) of the pairs in C4DET, while Pynguin and DPP generate a DET for 4.9% and 37.3% of the pairs, respectively. This promising result indicates that Mokav effectively leverages LLMs for differential testing. Mokav is the first LLM-based differential testing, and we provide a publicly available implementation for future research.
To summarize, we make the following contributions:
  • We propose a novel differential testing approach based on large-language models. It consists of iteratively giving execution feedback to an LLM to direct it towards generating a test input that exposes functional differences between two programs. The approach is implemented in Mokav, made publicly available for future research: https://github.com/ASSERT-KTH/Mokav
  • We introduce a large, carefully crafted dataset C4DET, collected from Codeforces competitions, to study differential testing in Python. C4DET contains 1,535 program pairs with guaranteed functional differences. We expect this dataset to facilitate future work on differential testing, with or without LLMs.
  • We present a large-scale, systematic evaluation of Mokav on QuixBugs and C4DET. Our results show that Mokav successfully generates a DET for 81.7% (1,255/1,535) of the pairs in C4DET, clearly outperforming the recent and strong state-of-the-art systems.
  • We conduct an ablation study of Mokav demonstrating the impact of each prompt component on its effectiveness. We also show that different LLMs (GPT-3.5, GPT-4o, and CodeLlama) can be successfully plugged into Mokav.
2. Problem Statement
In many software engineering tasks, developers face multiple versions of a program, having to reason about their functional difference or equivalence. For example, developers want to assess equivalence after refactoring. Another example is when developers need to ensure the functional difference of a mutant in mutation testing. A difference exposing test (DET) is a test input on which two given programs function differently at runtime (Taneja and Xie, 2008a), it is an existential, concrete proof of functional difference.
Definition 0 (Difference Exposing Test (DET)).
Let $P\colon I\to O$ and $Q\colon I\to O$ be two programs operating deterministically on the same input/output space, where $I$ is the input domain of $P$ and $Q$ and $O$ is the output domain. Then, det is a difference exposing test iff $det\in I\wedge P(det)\neq Q(det)$.
Generating a DETs is hard. The input space of the given programs can be excessively large, which makes it impossible to check the entire input space for finding a DET. Consequently, generating a DET requires an efficient exploration of the input space, based on guesses of potential functional differences between the given programs. We note that the execution of programs under analysis for each sample inputs can also take time (Pizzoleto et al., 2019).
Hence, an efficient DET generation tool must have two qualities: (1) leverage as much static information as possible for comprehending the behavior differences of given programs, and (2) perform a targeted input exploration to narrow down the input space before paying the price of execution. Both problems have been studied in recent years (VieiraB et al., 2022; Schoofs et al., 2022), but remain open.
Note that per Definition 2.1, a DET detects functional differences between P and Q, meaning they produce different outputs for a given input. In this work, we consider functional difference to be distinct from the execution difference. P and Q have an execution difference if, for a given input, there is a variable v used in both P and Q where the set of values v takes on in P is different from the set of values v takes on in Q. In essence, functional difference relates to external differences between P and Q that can be identified by an external observer; in contrast, execution difference pertains to internal differences between P and Q that are detected solely through monitoring the internal states of P and Q. The goal of this paper is to design a novel, effective approach for DET generation.
3. Approach
3.1. Overview
Figure 1. Overview of Mokav. Mokav exposes functional differences between two programs P and Q by feeding LLMs with execution feedback.
Figure 2. The initial prompt (pr0) used in Mokav.
1 Diff between P & Q:
2x = input().split()
3x.sort()
4for i in range(len(x)):
5 x[i] = int(x[i])
6x.sort()
7n = 100
8for j in range(2):
9 if ((x[j] + x[(j + 1)]) > x[(j + 2)]):
10 n = 300
11 elif ((x[j] + x[(j + 1)]) == x[(j + 2)]):
12 n = max(n, 200)
13 else:
14 n = max(n, 100)
15if (n == 300):
16 print(’TRIANGLE’)
17elif (n == 200):
18  print(’SIGMENT’)
19  print(’SEGMENT’)
20else:
21 print(’IMPOSSIBLE’)
22
23 Example test:
24Input:
25 "4 2 1 3"
26Output:
27 P: "TRIANGLE"
28 Q: "TRIANGLE"
29
30 Generated Difference-exposing test (DET):
31Input:
32 "5 2 1 3"
33Output:
34 P: "SIGMENT"
35 Q: "SEGMENT"
Listing 1: Two versions of a program that takes four numbers, sorts them, and checks if the first three or last three form a tirangle. An example test with identical output on P & Q and a DET generated by Mokav are also shown.
We design Mokav, a novel tool that uses an execution-driven approach for generating difference exposing tests (DETs) with large language models (LLMs). We implement it for generating DETs for Python programs.
Mokav takes an iterative approach to DET generation. It asks the LLM to generate a set of test inputs, gives feedback to the LLM regarding the execution data captured in the generated test inputs, and loops until a DET is generated or a maximum budget is reached.
Figure 1 overviews how Mokav works. Mokav takes as input two versions of a program (P and Q) as well as a sample test input, for which both versions produce the same output. We call this input an example test.
The example test is an essential component of Mokav for two reasons. First and foremost, Mokav utilizes the example test to extract valuable execution data. This execution data has two parts: (1) the output by P and Q, which is identical on the example test, and (2) execution differences between P and Q, if any, where an execution difference is a variable value that only occurs in one version, but does not lead to different outputs by P and Q. This detected execution difference is provided to the LLM to accurately steer it towards DET generation. Second, the example test provides valuable information for the LLM to understand the type and structure of the input space I of P and Q . Mokav’s outcome is a DET, an input data for which P and Q produce different outputs.
LABEL:lst:motivating_ex shows an example pair of programs given to Mokav and the DET it generates. In this example, P and Q are two real-world programs by a single user for a problem in the CodeForces [^2] https://codeforces.com/problemset/problem/6/A code competition. A correct program should sort the input numbers and check if the first three or the last three make up a triangle, a segment, or neither. The example test is “4 2 1 3”, for which both versions output “TRIANGLE”. Given P, Q, and the example test, Mokav generates a new input data: “5 2 1 3”. On this input, P’s output is “SIGMENT” with a misspelling, while Q correctly outputs “SEGMENT”. This means Mokav successfully generates a DET.
In the initial phase, Mokav creates an initial prompt based on P and Q, the example test, the execution data extracted while running the example test, and an LLM-generated description of the intention of P and Q. The description of P and Q clarifies the program intent in natural language, which is the language more familiar for many LLMs. Next, in the iteration phase, Mokav uses an LLM to generate new test inputs and iteratively gives execution based feedback to the LLM until it generates a DET. The generated DET is reported as the outcome of Mokav.
Our novel technique of generating DETs with LLMs employs two powerful features of LLMs. First, Mokav exploits the LLM’s ability to explore the input space for a given program. As previous work demonstrates, given appropriate information, LLMs can generate a high number of diverse and novel test inputs (Schäfer et al., 2023). This enables us to employ LLMs for exploring different parts of the input space of P and Q. Second, we leverage the LLM’s ability to understand the semantic differences between closely related programs. Recent studies show that advanced LLMs can accurately detect inconsistencies between programs and their descriptions (Li and Shin, 2024). This suggests that such LLMs can also accurately understand the differences between P and Q. We take advantage of this strength of LLMs to capture the functional differences between P and Q. Providing execution data to the LLM is a major part of Mokav’s design to improve LLMs’ understanding of P and Q.
3.2. Input
The input to Mokav is a tuple in the form of $(P,Q,T_{e})$. P and Q are a pair of versions of the program for which Mokav should generate a DET. The last input to Mokav, $T_{e}$, is an example test with the same corresponding output on both P and Q, which means it does not expose functional differences. This example test is used to capture execution differences between P and Q and also to show the type of the desired test input to the LLM (see Section 3.3). In LABEL:lst:motivating_ex, the example test hints to the model that the input contains four numbers, which makes the LLM less likely to generate invalid inputs with fewer or more numbers.
3.3. Mokav Initial Phase
In the initial phase, Mokav combines the inputs to generate an initial prompt, called pr0, that encompasses all the information required for starting the DET generation process.
First, Mokav generates two extra pieces of information: (1) execution data, which includes example test output and execution difference data, and (2) program descriptions. This information is obtained as follows:
(1) Example test output and execution difference: The runtime analyzer unit of Mokav executes the example test on P and Q to obtain the corresponding output and identify execution differences. Note that the output for the example test must be the same on both versions. If not, the generation is aborted. To identify the execution differences the runtime analyzer takes three steps as follows. First, it finds the variables that are common between P and Q. Next, it uses spotflow [^3] https://github.com/andrehora/spotflow to collect variable values during the execution of P and Q. Finally, the runtime analyzer identifies the first time a common variable gets a value in one of the versions, while that variable value never occurs in the other version. Such variable values are called unique variable values (Etemadi et al., 2023), and the runtime analyzer reports them as the detected execution difference. Note that depending on the example test, the runtime analyzer may be unable to detect a unique variable value for either of the versions. In this case, that part of the data is excluded from the prompt. The output for an example test and the detected execution differences clarify the behavior of the programs, which is crucial for DET generation.
(2) Program descriptions: Mokav prompts the LLM to generate natural language descriptions for P and Q. Recent studies show advanced LLMs are able to understand inconsistencies between programs and their natural language descriptions (Li and Shin, 2024). This suggests that adding the generated descriptions to pr0 can help the LLM to detect behavioral differences between P and Q. The prompt Mokav uses for description generation is a single line message “What is the intention of this code?”, followed by the program.
The pr0 follows the structure shown in Figure 2. The main part of the prompt consists of (1) the P and Q versions of the program, (2) the description of P and Q, (3) the example test and its corresponding output, (4) the execution difference detected while running the example test, and (5) text to specify the kind of response that is expected. For the latter, the expected format of the generated test is presented. Then, the prompt defines what is a difference exposing test. To ensure the LLM understands the definition, it is written in formal terms as well as natural language. The formal definition (“P(inputdate)!=Q(inputdata”) helps the model relate the DET generation task with the given P and Q programs.
From the initial prompt, the LLM generates a first candidate difference exposing test. Mokav executes both P and Q on the generated input to verify the presence of a functional difference. If that happens, Mokav generates a DET with the pr0 directly.
3.4. Mokav Iteration Phase
Mokav iteratively prompts the LLM until either it generates a DET or the maximum number of iterations is reached. At each iteration, after the model replies with a set of generated tests, the tests are executed on P and Q. If a test produces different outputs on P and Q, it is a valid DET, Mokav has succeeded, and we exit. Otherwise, Mokav tries again by requesting the LLM with an iteration prompt, called pr1. To create pr1, Mokav selects the first generated test and uses its runtime analyzer unit to produce execution based feedback for the selected test. This feedback contains the output of P and Q on the selected test and the execution difference detected during the execution of this test (see Section 3.3 for the execution difference detection description). Mokav gives feedback only on the first generated test to avoid exceeding the prompt size limit.
pr1 has three parts. The first part is the detected execution difference. If the runtime analyzer does not detect an execution difference, this part is excluded for pr1. The second part informs the LLM that P and Q produce the same output for the previously generated test, while we expect a test with different outputs on P and Q. The last part of the pr1 directly asks for another test.
At the end of the iteration phase, the result is either a generated test that exposes the functional difference or a failure message that the maximum number of iterations is reached.
3.5. Outcome of Mokav
The generated difference exposing test is presented to the user in the form of a unit test class. This test case imports P and Q as two functions fp and fq. Then a test case method is passes the generated DET to fp and fq and asserts that their outputs are not equal (for example, see the outcome section of Figure 1).
3.6. Implementation
Mokav is implemented in Python, and can be configured with different LLMs. By default, it supports OpenAI’s GPT models and CodeLLama-Instruct. It connects to the CodeLLama-Instruct API through the HuggingFace framework, which makes it easy to connect Mokav to many other open-source models as well.
Mokav uses three important values that can be configured for each execution: the maximum number of iterations, the number of samples requested from the LLM at each invocation, and the temperature of the LLM. The default configuration is 10 iterations, 10 samples, and temperature=1.
4. Experimental Methodology
4.1. Research Questions
In this paper, we study the following research questions:
  • RQ1 (effectiveness): How effective is Mokav for generating difference exposing tests? We run Mokav on QuixBugs (Lin et al., 2017) and a large dataset of pairs of programs with functional differences collected from Code4Bench (Majd et al., 2019). We assess the effectiveness of Mokav in DET generation and compare it against two baselines: (1) a state-of-the-art search-based test generation tool, called Pynguin (Lukasczyk and Fraser, 2022), and (2) DPP (Li et al., 2023).
  • RQ2 (ablation study): How does each component of Mokav impact its effectiveness? We evaluate whether and to what extent the different parts of Mokav impact its effectiveness.
  • RQ3 (impact of P and Q characteristics): How does the effectiveness of Mokav relate to the characteristics of the input programs? We check the effectiveness of Mokav in DET generation for different types of programs. More specifically, we check how the similarity between the two versions of the programs, the complexity of the test, and the length of the programs affect the effectiveness of Mokav.
4.2. Datasets
We employ two datasets in our experiments. First, QuixBugs (Lin et al., 2017) is a dataset of 40 pairs of programs in Python that implement common programming problems, such as the greatest common divisor. For each of the 40 pairs of programs in QuixBugs, one version is buggy and the other fixed. There is at least one input on which the buggy version either faces runtime error or has a return value different from the fixed version. This guarantees the functional difference between the two versions, with a significant functional overlap. This means that a DET is not easy to find. In this paper, we consider 32 program pairs from QuixBugs. For these pairs, QuixBugs contains a test that produces the same output on the buggy and fixed versions. This test should be used as the example test in Mokav’s workflow.
As a second dataset used in our evaluation, we create a curated dataset of pairs of programs, which we call C4DET. C4DET contains pairs of similar programs with guaranteed functional differences. We create C4DET by carefully collecting and curating submissions to the Codeforces code competition website. We use Code4Bench (Majd et al., 2019) to extract the required information for C4DET. Code4Bench contains problems, authors, submissions, and tests in Codeforces. Each submission is labeled with a verdict, which shows whether it passes the tests for the corresponding problem, gives a wrong answer, causes runtime error, causes timeout, or causes memory limit. C4DET consists of a subset of the programs and their corresponding problems and test cases in Code4Bench. These programs are selected as follows: for every author $a$ and every problem $pr$ in Codeforces, we add $\langle P,Q\rangle$ to C4DET iff they meet the following conditions:
  1. (1)
    $P$ and $Q$ are both Python submissions by $a$ for $pr$.
  2. (2)
    $P$ is the last submission by $a$ for $pr$ that gives a wrong answer for a test and $Q$ is the first submission by $a$ for $pr$ that passes all tests. This condition ensures two important conditions: (1) there is a functional difference between $P$ and $Q$, and (2) $P$ and $Q$ are closest submissions by $a$ for $pr$ that have a guaranteed functional difference.
  3. (3)
    There is at least one test input that produces the same output on $P$ and $Q$. We need this test in C4DET to use it as the example test.
  4. (4)
    $P$ and $Q$ are both shorter than 2,500 tokens when tokenized by the gpt-3.5 tokenizer. This condition is needed as the context size of LLMs is limited and excessively long programs cannot be processed by them.
  5. (5)
    The tests for $pr$ in Codeforces are shorter than 100 tokens when tokenized by the gpt-3.5 tokenizer. Again, this is needed as the context size of LLMs is limited, and long tests cannot be used as examples in the prompt.
  6. (6)
    There are at least two submissions in C4DET for $pr$ that pass all the tests. We add this condition to have at least two reference versions to be used for oracle assessment (Liu et al., 2024b).
The programs in Code4Bench are Python scripts that read inputs from stdin and print outputs to stdout. As many traditional test generation tools, such as Pynguin (Lukasczyk and Fraser, 2022), generate tests for functions, we transform all programs to their function-based versions in C4DET. LABEL:lst:function_based shows an example of this transformation. The source in Code4Bench is a Python script that reads a string (line 3) and prints the output (lines 6 and 8). The transformed version is a Python function that receives a list of arguments in *args (line 11) and creates a return_list (line 12) that is returned as the output of the function (line 21). The first input is read from args (line 15) and the outputs are added to the return_list at lines 18 and 20. This transformed version keeps the order of inputs and outputs as well as the functionality of the original source. At the same time, the transformed version is a function-based program suited for test generation without input/output.
Overall, C4DET contains 1,535 pairs of programs for 221. The longest pair of programs in C4DET has 4,242 tokens when tokenized by the gpt-3.5 tokenizer. Many advanced LLMs, such as gpt-3.5, have a 16,000 token limit. Thus, the maximum number of tokens in C4DET program pairs is low enough to be below this limit and leave space for other components of our prompt (see Figure 2). The median number of tokens in C4DET pairs of programs is 248. This is a reasonable number of tokens for a Python function. It is also close to the median number of tokens in Code4Bench, meaning that C4DET is representative of programs in online competitions, which are used in many software engineering research projects (Majd et al., 2019; Hendrycks et al., 2021).
1 Source of program P in Code4Bench:
2reg = re.compile(’(h)+(e)+(l)+(l)+(o)+’)
3 s1 = input()
4li = reg.findall(s1)
5if (not li):
6  print(’NO’)
7else:
8  print(’YES’)
9
10 Transformed version of P as a function-based program:
11 def P_func(*args):
12  return_list = []
13
14 reg = re.compile(’(h)+(e)+(l)+(l)+(o)+’)
15  s1 = args[0]
16 li = reg.findall(s1)
17 if (not li):
18  return_list.append(’NO’)
19 else:
20  return_list.append(’YES’)
21  return return_list
Listing 2: A program in C4DET and its function-based transformed version.
4.3. Baselines
We compare Mokav against two baselines: Pynguin (Lukasczyk and Fraser, 2022) and Differential Prompting (DP) (Li et al., 2023).
Pynguin is a traditional test generation tool that generates random unit tests for Python programs. Pynguin employs evolutionary algorithms (Panichella et al., 2017) to generate regression tests that maximize a fitness function, such as branch coverage. The generated tests call methods of a target module with correctly typed arguments and contain assertions about the methods’ return values. The assertions check the return values with simple types, i.e., int, str, bytes, bool, and None as well as collections of simple types. We consider Pynguin in our evaluation as it is a well-documented tool and recognized as state-of-the-art in test generation for Python (Li et al., 2023; Liu et al., 2024b).
The second baseline that we consider is DP (Li et al., 2023). DP is an LLM-based tool that gets a buggy program and generates fault-inducing tests for it. DP works in three steps. First, it asks the LLM to generate multiple fixed versions for the buggy program. Next, it requests the LLM to produce a set of test cases for the buggy version. Finally, it runs the generated tests on the buggy and fixed versions. If the generated tests produce different outputs on buggy and fixed programs, they are reported as fault-inducing tests.
Note that DP is different from Mokav, as it takes only a buggy program and produces fault-inducing tests for it. In contrast, Mokav takes two programs and generates difference exposing tests. Therefore, we cannot directly compare Mokav with DP. Instead, we isolate DP’s last step and call it DPP (standing for DP Plus). Similar to DP, DPP works in one iteration and generates all the requested tests in that one iteration. Moreover, it only provides one version of the program in the prompt and excludes the descriptions as well as the example test and its corresponding output and execution difference. The prompt is taken unmodified from the replication package of DP (Li et al., 2023).
4.4. Protocol RQ1 (effectiveness)
To answer RQ1, we run Mokav, Pynguin, and DPP on QuixBugs and C4DET. We use all the program pairs in C4DET for our experiments (see Section 4.2).
As Pynguin is designed for generating regression tests for one program, it cannot be directly used for DET generation. To address this challenge, we use Pynguin in three steps. For a given pair of programs $\langle P,Q\rangle$, we first run Pynguin 10 times on version P of the program, each time with a time budget of 1,000 seconds. Second, we run the generated tests on P and Q and record their outputs. Finally, if one of the generated tests produces different results on P and Q, we report it as a DET generated by Pynguin. Using this method, we can measure Pynguin’s effectiveness in DET generation.
We run Mokav for 10 iterations and generate 10 samples at each iteration. We also generate 100 samples in DPP’s single iteration. For this experiment, we select gpt-3.5-turbo-0125 as the LLM in use, as it provides a powerful and affordable model. To take advantage of gpt-3.5’s ability for generating diverse responses, we use temperature=1 in this experiment.
For each tool T, we run it on Quixbugs and C4DET and measure three metrics. First, we count the number of $\langle P,Q\rangle$ pairs for which T generates a DET, and call it #$\#Success\_Pairs$. This number is the main metric to compare the DET generation effectiveness of the considered tools. To better understand the effect of Mokav’s iterative approach, we measure this main metric for each of Mokav iterations. If Mokav generates DETs for more pairs after its first iteration, it indicates the effectiveness of the iterative approach. Second, we measure #$\#TT$, which is the total number of tests generated by T. #$\#TT$ indicates the scale of our study. Third, we measure $DET\_Pr$, the number of problems in C4DET that for at least one of their corresponding submission pairs $\langle P,Q\rangle$, T generates a DET. A tool that generates DETs for more problems is effective on a more diverse set of programs, indicating strong external validity.
This experiment is conducted on a machine with 128 AMD EPYC 7742 64-Core processors, running at 2.2GHz and having eight 16GB DDR4 RAMs with 3.3GHz speed.
Table 1. Effectiveness of Mokav compared with related work. QuixBugs dataset contains 32 pairs of programs, and C4DET contains 1,535 pairs of programs for 221 problems. #Success_pair is the number of pairs for which a DET is generated, #TT is the number of all generated tests, and #DET_Pr is the number of problems a DET is generated for at least one of the corresponding pairs.
QuixBugs C4DET
Tool #Success_Pair #TT #DET_Pr #Success_Pair #TT #DET_Pr
Pynguin 50.0% (16/32) 1,176 50.0% (16/32) 4.9% (76/1,535) 27,487 18.5% (41/221)
DPP 50.0% (16/32) 3,200 50.0% (16/32) 37.3% (573/1,535) 153,500 70.5% (156/221)
Mokav (iter 1) 93.6% (30/32) 320 93.6% (30/32) 58.9% (905/1,535) 15,350 87.7% (194/221)
Mokav (iter 4) 100.0% (32/32) 340 100.0% (32/32) 76.4% (1,173/1,535) 30,260 90.9% (201/221)
Mokav (iter 7) 100.0% (32/32) 340 100.0% (32/32) 80.3% (1,234/1,535) 40,240 92.7% (205/221)
Mokav 100.0% (32/32) 340 100.0% (32/32) 81.7% (1,255/1,535) 48,930 94.1% (208/221)
4.5. Protocol for RQ2 (ablation study)
To answer RQ2, we run Mokav on C4DET and count #$\#Success\_Pairs$, explained in Section 4.4, with various settings as follows:
  • To measure the effect of descriptions, we run Mokav without providing the program descriptions in the prompt.
  • To measure the effect of randomness in LLM’s responses, we run Mokav with temperature=0. This leads to more deterministic responses, meaning we generate only one sample at each iteration for this setting.
  • To measure the effect of LLM’s accuracy and check if Mokav is well designed to work with various LLMs, we also run Mokav using two other LLMs. First, we use
    CodeLlama-instruct-7b as an open-source LLM. To keep the resource demand of this experiment manageable, we run this experiment with minimum temperature and generate one sample at each iteration. Second, we also run Mokav with gpt-4o-2024-05-13, as one of the most powerful existing LLMs. We use GPT-4o with two configurations: (1) the same configuration as for CodeLlama, to compare all models with each other, and (2) the default configuration used by Mokav, to assess Mokav’s effectiveness when using the most advanced existing LLM.
  • To measure the effect of example test, we run Mokav without providing the example test and its corresponding output and execution difference in the pr0 and pr1 prompts.
  • To measure the effect of execution data, we run Mokav without providing the output and execution difference in the pr0 and pr1 prompts.
4.6. Protocol for RQ3 (impact of P and Q characteristics)
To answer RQ3, we assess the relation between the effectiveness of Mokav and three different characteristics of the input to the tool: the number of tokens in P and Q, the Levenshtein distance between P and Q, and the number of tokens in the longest test for program pair in C4DET. Note that C4DET contains multiple tests for each problem in C4DET. We consider the longest test in this experiment as it suggests how complex the input to the program pair can be.
[…]
To evaluate the relation between DET generation effectiveness and the two other input characteristics (distance between P and Q, and number of tokens in the example test), we follow a similar strategy. The only difference is the metric according to which we sort the $\langle P,Q\rangle$ pairs and split the data into ten subsets. A significant correlation between $SSuccess$ and each of the three considered metrics provides valuable information regarding where we should expect Mokav to work in practice. It can also inform us what are the fundamental reasons why LLMs can generate DETs in some cases and fail to do so in other cases.
5. Experimental Results
5.1. Results for RQ1 (effectiveness)
Figure 3. Effectiveness over iteration. Mokav drives the LLM towards finding difference-exposing tests.
1 The diff between P & Q:
2s = input()
3s1 = [i for i in range(len(s)) if ((s[i]==’A’) or (s[i]==’E’) or (s[i]==’I’) or (s[i]==’O’) or (s[i]==’U’) or (s[i]==’Y’))]
4d = 0
5for i in range((len(s1) - 1)):
6 if ((s1[(i + 1)] - s1[i]) > d):
7 d = (s1[(i + 1)] - s1[i])
8if ((d == 0) and (len(s1) != 0)):
9 print(max((s1[0] + 1), (len(s) - s1[(- 1)])))
10elif (len(s1) == 0):
11 print((len(s) + 1))
12else:
13  d = max(d, (len(s) - s1[(- 1)]))
14  d = max(d, (len(s) - s1[(- 1)]), (s1[0] + 1))
15 print(d)
16
17 DET generated by Mokav:
18Input:
19 "BYAIAUOIEOAA"
20Output:
21 P: 1
22 Q: 2
Listing 3: A pair of programs for which only Mokav is able to generates a complex DET, after several iterations.
Table 1 compares the effectiveness of Mokav on QuixBugs and C4DET with two baselines (Pyguin and DPP). For each row, the first column is the name of the tool. The table presents Mokav’s effectiveness with 1, 4, and 7 iterations, and its default setting with 10 iterations for the last row. For each dataset, we measure three metrics: the number of program pairs with DETs (“#Success_Pair”), the total number of generated tests (“#TT”), and the number of problems that the tool generates a DET for at least one of its corresponding program pairs (“#DET_Pr”), respectively.
Table 2. Ablation study about the effectiveness of different Mokav components.
Config ID LLM Iterations Samples Temperature Description Ex_Test Exec_Data #Success_Pairs
C1 (Mokav) GPT-3.5 10 10 1 81.7% (1,255/1,535)
C2 GPT-3.5 10 10 1 80.4% (1,235/1,535)
C3 GPT-3.5 10 10 1 26.3% (404/1,535)
C4 GPT-3.5 10 10 1 79.4% (1,220/1,535)
C5 GPT-3.5 10 1 0 51.7% (795/1,535)
C6 CodeLlama 10 1 0 11.4% (175/1,535)
C7 GPT-4o 10 1 0 73.9% (1,135/1,535)
C8 GPT-4o 10 10 1 85.5% (1,313/1,535)
According to Table 1, Mokav generates a DET for 58.9% (905/1,535) of C4DET pairs at its first iteration and for 81.7% (1,535/1,535) of pairs at its last iteration, which is the highest performance ever reported. First, this shows that the iterative approach significantly improves the effectiveness of Mokav. This is also illustrated in Figure 3. Second, this result shows that Mokav outperforms both strong baselines Pynguin and DPP, which generate DETs for 4.9% and 37.3% of program pairs, respectively. Our manual investigation reveals that Mokav outperforms Pynguin and DPP mainly because it better understands the type of arguments the programs take. This confirms the importance of using an example test in the prompt to hint the LLM regarding the type and structure of inputs.
We notice that Mokav, due to its iterative and execution-driven feedback, better understands the semantic differences between P and Q. LABEL:lst:outperformance_ex shows an example where Mokav generates a DET at its final iteration, while Pynguin, and DPP fail to do so. P and Q are supposed to compute the distance between certain characters in a string and the beginning and end of a given string, and output the longest distance. P wrongly ignores the distance between the beginning of the string and the first appearance of the considered characters (line 13). Mokav successfully generates “BYAIAUOIEOAA” as a DET for this pair. The DET is only generated at the last iteration of Mokav. This indicates that the iterative and execution-driven approach of Mokav enables it to better understand P and Q’s semantics compared to DPP and Pynguin.
The next column (“#TT”) shows the total number of tests generated by each tool. DPP generates 153,500 tests for 1,535 pairs in C4DET, because it asks for a fixed number of 100 tests for each pair. Mokav generates batches of 10 tests at each iteration only if the previous iterations fail to yield a DET. Consequently, Mokav generates only 15,350 tests in its first iteration and this number gradually increases to 48,930 at its 10th iteration. This shows that Mokav’s iterative approach also contributes to its cost efficiency by generating new tests step-by-step and only when needed.
The “#DET_Pr” column of Table 1 presents the number of problems for which the tool generates a DET for at least one of their corresponding program pairs. Recall that the program pairs in C4DET are submissions to answer online competition problems. A problem may have many corresponding program pairs in C4DET (see Section 4.2). In QuixBugs, there is exactly one pair per problem; therefore, #Success_Pair and #DET_Pr are the same for QuixBugs. Mokav outperforms both baselines in terms of #$\#DET\_Pr$ by generating a DET for 94.1% (208/221) of the problems. This shows that Mokav’s effectiveness is not limited to small and specific types of programs; it can generate DETs for a diverse set of programs.
All the major takeaways from the results on C4DET also hold for QuixBugs. Most importantly, Mokav outperforms both Pynguin and DPP even at its first iteration. Mokav generates a DET for 100% (32/32) of the program pairs in QuixBugs. This demonstrates the perfect effectiveness of Mokav on this widely-used dataset.
{mdframed}
Answer to RQ1: How effective is Mokav for generating difference exposing tests?
We evaluate DET generation with Mokav, Pynguin, and DPP on 32 program pairs from QuixBugs and 1,535 program pairs from C4DET. The results clearly show that Mokav outperforms both baselines on both datasets. On QuixBugs, Mokav achieves a perfect score by generating DETs for 100% (31/32) of cases. On C4DET, Mokav produces DETs for 81.7% (1,255/1,535) of the program pairs. Our results also demonstrate that Mokav’s effectiveness improves over iterations, starting at 58.9% and improving up to 81.7%. To our knowledge, Mokav is the best performing, state-of-the-art approach and tool for generating DETs.
1 The diff between P and Q:
2# P and Q get two integers (n and k) and a string (a1) as input and return a list of strings as output.
3...
4  - if ((1 <= n <= 50) and (1 <= k <= 50)):
5  + if ((1 <= n < 50) and (1 <= k <= 50)):
6...
7
8 The DET generated by Mokav:
9{n=50, k=1, a1=’BBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBG’ }
10
11 The output produced by P and Q on the generated DET:
12P output: []
13Q output: [’BBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBGB’]
Listing 4: Code diff between a pair of programs for problem 226B. A DET for this pair of programs should be an input string that contains exactly 50 characters. GPT-4o is the only model that generates a DET for this program pair.
5.2. Results for RQ2 (ablation study)
Table 2 presents the results of our ablation study. It compares the effectiveness of different Mokav configurations in terms of their #Success_Pairs. Eight configurations, C1-C8, are considered, with C1 being the default setup of Mokav. C2, C3, and C4 represent configurations where the description, the example test, and the execution data are excluded from the prompt, respectively. Note that in C3, as the example test is excluded, the execution data that depends on the presence of an example test is excluded as well. In C5, temperature=0 is used and only the top one sample is requested from the model at each iteration. C6 and C7 are similar to C5, but use different LLMs, GPT-4o and CodeLlama, respectively. Finally, C8 has the same configuration as C1, but uses GPT-4o as the LLM. The last column of Table 2 shows the #Success_Pairs, the number of pairs in C4DET for which each configuration generates a DET.
C2, C3, and C4 show that the presence of each component of the prompt contributes to the effectiveness of Mokav. C3 results show that removing the example test and its corresponding execution data has the highest impact on effectiveness by decreasing it from 81.7% to 26.3%. Our manual investigation reveals the main reason for this decline: by removing the example test, the LLM cannot recognize the type and structure of the input to be generated for the program pairs. For example, the example test hints to the LLM if the inputs should be strings or integers, what should be the content of string inputs, how many inputs are required, etc. In the absence of the one-shot example test, all this information is missing from the prompt and this negatively affects the effectiveness. Overall, the example test has the most significant impact.
The effectiveness of C5 is 51.7%, versus 81.7% for C1. This shows the large impact of using a high temperature and sample size. By using a high temperature and sample size, we can better exploit the LLM’s capability to explore the input space of program pairs.
C6, C7, and C8 use different LLMs. The table shows that the model in use influences effectiveness. CodeLlama underperforms with an effectiveness of only 11.4% (C6 $\ll$ C5). Expectedly, the very recent GPT-4o (released in May 2024) surpasses GPT-3.5 (C7 $\gg$ C5).
Finally, C8 shows results for the last experiment, where we run Mokav with default settings similar to C1 using GPT-4o, released while we were preparing this manuscript. The results show that the effectiveness is improved to 85.5%, the highest among all configurations. This improvement compared to C1 shows that using a stronger LLM that better understands the semantics of programs directly impacts effectiveness. We note that C8 has an 11.6% improvement over C7 (85.5% vs. 73.9%). This significant improvement reaffirms the soundness of the default setting used in Mokav: a high sample size (10) and high temperature (10) lead to generating a diverse set of useful test cases.
We now discuss an interesting case in LABEL:lst:gpt4-ex, which shows the code diff between a pair of programs, and input and output. The pair comes from author “60,724” for problem 226B on Codeforces. Only GPT-4.o succeeds in generating a valid DET, with exactly 50 characters as input (“BBBBBBBBBBBBBBBBBBBBBBBBBBBBBBB
BBBBBBBBBBBBBBBBBBG”). This shows that the increasingly more powerful LLMs can be very valuable for generating DETs.
On the negative side of results, there are 11.5% (178/1,535) of pairs that no config generates a DET for. LABEL:lst:no-det is one of these pairs. Per our manual analysis, in most of these pairs, similar to LABEL:lst:no-det, the LLM needs to perform precise mathematical computations to generate a DET. This reveals a core limitation of LLMs that affects their performance on code related tasks.
{mdframed}
Answer to RQ2: How does each component of Mokav impact its effectiveness?
Our comparison of 8 different configurations of Mokav shows that using Mokav’s default setup and GPT-4o surpasses all other configurations with an effectiveness of 85.5% (1,313/1,535). Among components in Mokav’s approach, the example test given tothe LLM plays the most important role. Using a high temperature and sample size effectively leverages the LLM’s ability to explore various parts of the input space.
5.3. Results for RQ3 (impact of P and Q characteristics)
Table 3. The correlation between Mokav’s effectiveness and P & Q characteristics.
SRC_Tok Test_Tok Lev_Diff
Med SSuccess Med SSuccess Med SSuccess
S1 74 0.96 1 0.89 0.44 0.92
S2 115 0.88 2 0.84 0.68 0.90
S3 153 0.84 3 0.79 0.80 0.86
S4 187 0.86 4 0.77 0.87 0.76
S5 220 0.77 6 0.77 0.91 0.76
S6 263 0.79 9 0.83 0.94 0.79
S7 312 0.77 13 0.78 0.96 0.77
S8 384 0.81 21 0.68 0.99 0.80
S9 496 0.75 55 0.84 0.99 0.79
S10 809.5 0.72 82 0.92 0.99 0.78
Cor -0.89 -0.01 -0.45
p-val(<0.05) 0.0004 (✓) 0.97 (✗) 0.18 (✗)
1 The diff between P and Q:
2n, k = map(int, input().split())
3 s = int(n / k)
4 s = int(n // k)
5if s   2 != 0:
6...
Listing 5: A pair for which no config generates a DET.
Table 3 shows the results of the RQ3 experiment. In this table, S1-S10 represent the subsets of program pairs after split according to each considered program pair characteristics. The table has three main vertical sections representing each of those three characteristics: “SRC_Tok”, the number of tokens in a program pair; “Test_Tok”, the number of tokens in the longest test in C4DET for a program pair; and “Lev_Diff”, the Levenshtein distance between a pair of programs. For each characteristic, the table shows the median value for each subset (“Med”) and the ratio of pairs in each subset for which Mokav successfully generates a DET (“SSuccess”). For example, consider S1. The median number of tokens in the 10% shortest pairs is 74. Mokav generates a DET for 96% of these program pairs. The median number of tokens in the test for 10% of program pairs that have the shortest test is 1. Mokav generates a DET for 89% of these program pairs. Finally, the median Levenshtein distance for the 10% of program pairs with the shortest distance is 0.44. For 92% of these program pairs, Mokav successfully generates a DET.
The last two rows of Table 3 show the correlation between the subset index and each of the considered characteristics. There is a strong negative correlation of -0.89 between the number of tokens in source code and the success rate of Mokav, meaning that the longer the input programs, the less likely Mokav can generate a valid DET. This correlation is statistically significant as its corresponding p-value is 0.0004, much lower than the typical 0.05 threshold. Overall, Table 3 shows that the number of tokens in program pairs is the only characteristic with a statistically significant correlation with Mokav’s success rate. This indicates that to apply Mokav on multi-function programs, we would have to employ powerful LLMs with large context windows, and delicately provide them with runtime hints regarding program semantics.
The absence of a significant correlation between the length of the longest test in C4DET for a program pair and Mokav’s success rate demonstrates the following: the size of the inputs for a program pair does not directly tell how challenging it is to find the corner cases that expose differences. We observe no significant correlation between the program pair’s distance and Mokav’s success rate. At first glance, it is easier to expose the difference between two programs if they are textually distant; therefore, we expect to see a positive and significant correlation between Mokav’s success rate and the distance between pairs. On the other hand, we notice that in C4DET, program pairs close to each other tend to be small and easier to understand. This means we expect the LLM can better understand program pairs with small distances and thus generate a valid DET for them. Consequently, the distance between program pairs impacts Mokav’s success rate from two opposite directions. This leads to the absence of a significant correlation between Mokav’s effectiveness and distance between P&Q.
{mdframed}
Answer to RQ3: How does the effectiveness of Mokav relate to the characteristics of the input programs?
Our experiments show a significant correlation between the length of program pairs and Mokav’s success rate in generating a DET for them. There is strong statistical evidence that LLMs better capture the semantics of small program pairs to find corner cases that expose their differences. We do not find a significant correlation between Mokav’s effectiveness and the distance between program pairs.
6. Threats to Validity
Construct Validity: The main threat to the construct validity of this study is data leakage in LLMs (Al-Kaswan et al., 2024). As LLMs are trained on a large corpus of open-source data, they may have some program pairs of our datasets in their training data. This may threaten the generalizability of our experimental conclusions. Despite this concern, we note open-source repositories usually do not put difference-exposing tests next to programs. This means there is no direct connection between programs and DETs in LLMs’ training data. Consequently, the data leakage problem is less significant in DET generation, the goal of Mokav, compared to more traditional tasks, such as code generation (Du et al., 2024) and program repair (Jiang et al., 2023). Also, to further alleviate the data leakage concern, we assess the effectiveness of Mokav on two different datasets, i.e., QuixBugs (Lin et al., 2017) and C4DET.
Internal Validity: Mokav uses a high temperature for the LLM in its default configuration. This leads to randomness in its responses, a threat to the internal validity of our experimental results. To address this threat, we repeat our main experiment of generating DETs for program pairs in C4DET five times. The results show that in all five instances, the effectiveness of Mokav is between 80.1% and 81.7%. This indicates that the main conclusions of our study are robust and only slightly affected by the randomness of LLMs.
7. Related Work
Test Generation with LLMs LLMs are being applied on many software engineering tasks (Fan et al., 2023), and test generation is one of the major targets (Wang et al., 2024; Siddiq et al., 2023; Xie et al., 2023; Tang et al., 2024; Lemieux et al., 2023; Tufano et al., 2020; Altmayer Pizzorno and Berger, 2024; Chen et al., 2022; Takerngsaksiri et al., [n. d.]; Plein et al., 2023; Lahiri et al., 2022; Vikram et al., 2023; Liu et al., 2024a; Nie et al., 2023; Ryan et al., 2024; Dakhel et al., 2024; Pizzorno and Berger, 2024; Alshahwan et al., 2024; Steenhoek et al., 2023). A test consists of multiple parts: an input to a program under test (PUT), a test setup that moves the PUT into a specific state, and an oracle that asserts the PUT correctly produces the expected outputs. LLMs are used to generate all these parts of tests.
Deng et al. (Deng et al., 2023) introduce TitanFuzz. This tool employs LLMs to generate human-like code that tests the API of deep learning (DL) libraries. TitanFuzz uses LLMs to generate and mutate input DL programs for fuzzing. Xia et al. (Xia et al., 2024) propose Fuzz4All, an LLM-based fuzzer that generates input test programs for compilers. Fuzz4All iteratively prompts the LLM to create new test inputs. The authors show that Fuzz4All effectively generates test inputs for six different programming languages. While TitanFuzz and Fuzz4All generate input programs for a single system, Mokav generates test inputs that differentiate two given programs.
LLMs have also been used for oracle generation (Watson et al., 2020; Tufano et al., 2022; Hossain and Dwyer, 2024; Li and Yuan, 2024). Dinella et al. propose TOGA (Dinella et al., 2022). TOGA uses a fine-tuned version of CodeBERT (Feng et al., 2020) to generate assertion candidates based on the PUT and test prefix and rank these candidate assertions. Liu et al. (Liu et al., 2023b) show limitations of TOGA: its unrealistic assumption of the availability of a correct focal method, its ignorance of important evaluation metrics, such as precision, and the lack of a straightforward baseline. Zhang et al. (Zhang et al., 2024) generate tests in which the oracle expects an exception. In contrast to these works, Mokav focuses on input generation instead of oracle generation.
Li et al. (Li et al., 2023) introduce differential prompting, an LLM-based tool that generates fault-inducing tests. This tool takes a buggy program. It first uses the LLM to generate an intention description for the program. Then, it generates a fixed version of the program. Next, it employs the LLM to generate a list of test inputs. Finally, a test input that leads to different results on the buggy and fixed versions is labeled as fault-inducing. AID (Liu et al., 2024b) is another tool that follows the same procedure, while it uses the LLM to generate a test generator, instead of test inputs. These two tools are different from Mokav as they generate fault-inducing tests for a given buggy program. In contrast, Mokav directly generates difference exposing tests for two given programs.
Liu et al. (Liu et al., 2024c) use LLMs to amplify existing tests in the HumanEval dataset (Chen et al., 2021). They generate new tests that expose differences between the dataset programs and incorrect programs generated for the code documentation. Their test generation method starts from test seeds generated by LLMs and then improves them by type-aware mutation. This method is different from Mokav, which uses an execution-driven iterative approach to improve initial tests generated by the LLM.
Generating Tests for Python Generating tests for dynamically typed programming languages is challenging (Selakovic et al., 2018). Auger (Laffra, 2024) is one of the first tools to address this task for Python. Auger uses Python tracer to collect the values with which target functions are called and the values that target functions return. The collected data is then used to generate unit tests.
The state-of-the-art tool for Python test generation is Pynguin (Lukasczyk and Fraser, 2022; Lukasczyk et al., 2020; Erni et al., 2024), which uses evolutionary algorithms to generate unit tests with high coverage. It uses type annotations and static analysis to infer the type of function parameters. Lukasczyk et al. (Lukasczyk et al., 2023) conduct a study to improve Pynguin’s initial type inference strategy and find a more suitable evolutionary algorithm: DynaMOSA (Panichella et al., 2017).
Schoofs et al. (Schoofs et al., 2022) introduce AmPyfier, which amplifies existing Python test suites. AmPyfier runs existing tests on the PUT and collects values during execution. These collected values are then added to the existing test as new assertions. AmPyfier also amplifies the inputs to PUT iteratively. To make the amplification efficient, at each iteration, AmPyfier only keeps the assertion and input amplifications that increase mutation or code coverage score.
In contrast with previous Python test generation tools, Mokav focuses on difference-exposing test generation.
Differential Testing Differential testing is a well-established technique for detecting anomalies and defects in programs (McKeeman, 1998; Evans and Savoia, 2007; Taneja and Xie, 2008b; Nilizadeh et al., 2019; Rutledge and Orso, 2022; Ye et al., 2023). It is employed to identify issues in many types of programs, such as robotic vehicle programs (Kim et al., 2023), game agents (Castellano et al., 2022), spreadsheet applications (Almasi et al., 2018), certificate parsers (Chen et al., 2023), and model counters (Usman et al., 2020). Research has shown the usefulness of differential testing in large-scale industrial software projects(Gulzar et al., 2019; Jarman et al., 2020a, b).
Silva et al. (Da Silva et al., 2020) study the feasibility of detecting semantic conflicts at merge time with differential testing. They use EvoSuite (Fraser and Arcuri, 2011) and Randoop (Pacheco and Ernst, 2007) to generate tests before a branch is merged. They show the generated tests can reveal semantic conflicts between the base and merging branches.
Recently, differential testing has also been used for testing deep learning programs and networks (Zhang et al., 2021). Guo et al. (Guo et al., 2018) propose Dlfuzz, a tool that generates rare inputs for DL networks. Dlfuzz uses a mutation algorithm to generate new inputs that activate previously unactivated neurons. Liu et al. (Liu et al., 2023a) propose a novel technique, called Gandalf, that generates inputs to detect inconsistencies between DL libraries.
Mokav is the first LLM-based differential testing tool, which uses an execution-driven and iterative prompting approach.
8. Conclusion
In this paper, we introduce Mokav, a novel tool that generates difference exposing tests (DET) employing LLMs with an execution-driven iterative approach. Mokav takes two versions of a program (P & Q) and an example test that produces the same output on P & Q. Mokav’s outcome is a test input that produces different results on P & Q. Mokav works with advanced prompting of an LLM: the prompt consists of P & Q, the example test, the output produced by the example test, and monitoring data collected during the execution of the example test. We evaluate Mokav on QuixBugs and a novel benchmark called C4DET, both in Python. Our experiments reveal that Mokav is effective and generates a DET for 81.7% (1,255/1,535) of program pairs in C4DET and 100% (32/32) of program pairs in QuixBugs. Mokav significantly outperforms Pynguin (Lukasczyk and Fraser, 2022) and DPP (Li et al., 2023), the closest, state-of-the-art test generation tools for the considered task.